For each pair of numbers find integers π‘₯ and 𝑦 such that π‘Žπ‘₯ + 𝑏𝑦 =gcd (π‘Ž,𝑏)


a) π‘Ž= 91, 𝑏=10



1
Expert's answer
2021-11-15T18:32:35-0500

For the pair "\ud835\udc4e= 91, \ud835\udc4f=10" of numbers let us find integers "x" and "y" such that "\ud835\udc4e\ud835\udc65 + \ud835\udc4f\ud835\udc66 =gcd (\ud835\udc4e,\ud835\udc4f)."

Taking into account that "91=10\cdot9+1" and "10=10\cdot 1+0," we conclude that "gcd (91,10)=1=91-10\cdot 9=91\cdot 1+10\cdot(-9)." Consequently, "x=1,\ y=-9."